# Detecting Machine-Generated Music with Explainability – A Challenge and Early Benchmarks

### v1

@@ IV-B Quantitative Analysis
2 29.6$\pm$7.9 48.6$\pm$4.4 - - -
3 10.2$\pm$2.2 55.6$\pm$5.2 0.15 65.5 14.4
4 3.42$\pm$0.4 76.0$\pm$3.0 0.004 66.5 36.7
5 0.43$\pm$0.3 80.0$\pm$2.2 0.01 87.4 5.3
V Conclusion
@@
In this work, we benchmarked the performance of seven models spanning traditional machine learning, deep neural networks, and Transformer-based methods on the MGMD task, providing a robust set of baselines for future research. Additionally, we explored multimodal approaches, demonstrating their potential in advancing music feature analysis by incorporating complementary modalities such as lyrics.
@@
To assess model robustness and generalisability, we conducted out-of-domain testing and utilised XAI techniques to analyse whether the models learnt intrinsic features. Furthermore, we proposed a novel ensemble XAI strategy to interpret model decisions effectively. Our findings reveal that existing models often fail to leverage fundamental musicological features, such as stops and structural elements, indicating a gap in their ability to capture domain-specific knowledge. This highlights the need for future research to integrate domain-specific knowledge into model architectures, paving the way for more interpretable and robust AI systems in music analysis.

### v2

@@ Introduction
[…]
Table 7: Fidelity experiments of combined overlapping regions from multiple XAI techniques
Discussion
This study establishes a comprehensive systematic evaluation for the task of MGMD. A diverse array of architectures are evaluated ranging from traditional CNNs to emerging State Space Models. Our extensive quantitative experiments confirm that while binary classification is achievable with high accuracy using deep neural networks within specific domains. By systematically assessing these models, we validate the feasibility of deploying automated detection systems whilst highlighting the performance degradation observed when models confront unseen generative algorithms or hybrid datasets. Furthermore, the integration of multimodal features and XAI in our systematic evaluation. Beyond binary classification, this study has remarkable implications for real-world deployment, such as automated copyright management and content tagging on streaming platforms. A robust detector can protect human artists by filtering massive volumes of AI-generated uploads.
Based on the systematic evaluation results, several practical observations can be drawn to guide the future research. Regarding model choice, CNN-based architectures should be favored when training data and representations are fixed, as they consistently achieve strong performance across both in-domain and out-of-domain evaluations, whereas Transformer-style models (such as ViT) do not show a clear advantage and in some cases underperform despite higher computational cost. Concerning input representations, all models are evaluated using the same spectrogram configuration, and the strong performance achieved by multiple CNN variants indicates that standard spectrogram settings are already sufficient to support competitive MGMD performance, without requiring task-specific feature redesign. However, in the multimodal setting, a simple early-fusion strategy based on concatenating frozen audio and text embeddings yields consistent improvements over unimodal baselines, suggesting that lightweight fusion mechanisms constitute an effective and practical starting point, while more complex multimodal modeling can be explored in future work.
From a musicological perspective, our XAI analysis points to a probable divergence between model behaviour and human auditory perception. As observed in the specific case study (Fig. 3), the model appeared to penalise a ’musical break’ (silence), potentially treating it as a spectral defect. In contrast, human listeners perceive such pauses as structural rhythmic elements. Although establishing this pattern as a systematic bias requires larger-scale expert annotation, our findings provide a compelling hypothesis: that current models heavily rely on low-level spectral artefacts rather than high-level musical semantics. Future MGM detection should aim to validate and bridge this gap by incorporating music-theoretic features such as beat tracking and structural segmentation to align model judgements closer to human aesthetic perception.
Note that all analyses based on the evaluation results are conducted within a general-purpose audio processing framework, and we adopt a standard binary supervised learning setting commonly used in audio deepfake detection. The results indicate that a substantial degree of shared knowledge exists between audio and music deepfake detection. Therefore, as this work aims to provide an early systematic evaluation, we do not introduce novel methodologies, loss functions, or model architectures. Nevertheless, the XAI techniques reveal important regions relevant to the MGMD task, and, together with the musicological analysis, they suggest a promising direction for future research, as illustrated in Figure 5. Specifically, acoustic features emphasising particular frequency regions, raw deep learning-based audio representations, semantic information (e.g. derived from lyrics), and musicological features such as rhythm could be carefully extracted and more effectively fused using specially designed mechanisms, potentially inspired by cross-attention. These fused representations could then be further processed in a post-feature stage for classification.
Figure 5: Proposed MGMD pipeline. The music is first processed to extract diverse features, including acoustic, semantic, musicological, and deep learning–based representations. These features are subsequently processed and fused, and the fused representations are ultimately used for classification.
Also, we acknowledge several limitations. First, We relied on Mel-spectrograms and basic early fusion for multimodal tasks. While sufficient for establishing baselines, as it is currently the most widely adopted representation in modern audio models, including popular systems such as Whisper [32] and recent audio LLMs [30], this approach may overlook temporal structural features that are crucial in musicology. We focus on FakeMusicCaps and M6, as they are among the few publicly available music deepfake datasets, whereas other existing datasets are not publicly accessible or have legal constraints. Moreover, inviting professional musicians to judge based on musicology could greatly add more insight to future feature engineering. In addition, as noted in our qualitative analysis, current XAI methods primarily visualise spatial attention rather than providing semantic reasoning. They highlight spectral regions but do not explicitly explain decision logic based on high-level musical attributes like rhythm or harmony.
Conclusion
@@
In this work, we systematic evaluate the performance of ten models spanning traditional machine learning, deep neural networks, and Transformer-based methods on the MGMD task, providing a robust set of baselines for future research. Additionally, we explored multimodal approaches, demonstrating their potential in advancing music feature analysis by incorporating complementary modalities such as lyrics.
@@
To assess model robustness and generalisability, we conducted out-of-domain testing and utilised XAI techniques to analyse whether the models learnt intrinsic features. Furthermore, we proposed a novel ensemble XAI strategy to interpret model decisions effectively. Our findings reveal that existing models often fail to leverage fundamental musicological features, such as stops and structural elements, indicating a gap in their ability to capture domain-specific knowledge. This highlights the need for future research to integrate domain-specific knowledge into model architectures, paving the way for more interpretable and robust AI systems in music analysis. Additionally, MGMD based on alternative music representations, such as symbolic or MIDI-based formats, could be explored in future work once more suitable datasets become available.
Data Availability
All data generated or analysed during this study are included in FakeMusicCaps: https://paperswithcode.com/dataset/musiccaps, and M6 [24] published article.
Funding
This research was partially supported and funded by the Munich Center for Machine Learning and the Munich Data Science Institute.
Author contributions statement
Yupei Li: lead the project, write manuscripts, do experiments; Qiyang Sun: write manuscripts, do experiments; Hanqian Li: do experiments; Lucia Specia: supervision, check manuscripts, provide GPUs; Björn W. Schuller: major supervision, re-write manuscripts, funding